What is the vertex of g(x)=x^2+9x+18
Answer by stanbon(75887)

What is the vertex of g(x)=x^2+9x+18
---
Vertex occurs at x = -b/(2a) = -9/(2*1) = -4.5
-----
g(-4.5) = (-4.5)^2+9(-4.5)+18 = -2.25
----------------------------------------
Vertex(-4.5 , -2.25)
========================
Cheers,
Stan H.
==================